By selling a watch for rs.1275, javed lost 15%. At what price should he sell to kae a profit of 10%
Found 2 solutions by Maths68, mananth:
Answer by Maths68(1474)


Let
Original Price of Watch = x
15% of x = x*15% = x*0.15
Given
Selling pirce = 1275
Original Price - 15% of original pirce = Selling Price
x-0.15x=1275
0.85x=1275
0.85x/0.85=1275/0.85
x=1500
Original Price of Watch = Rs. 1500
10% of 1500 = 10/100 * 1500 = 1/10 * 1500 = 1500/10 = 150
Selling price at 10% profit = 1500 + 150 = 1650


He sould sell watch to kae for Rs. 1650 to get a profit of 10%.

Answer by mananth(16949)

SP= 1275
CP= x
loss 15
CP=85
CP=100
85/100=1275/x
85x=127500
x=127500/85
x= 1500